OpenStudy (anonymous):

5 - (6k + 1) = 2[(5k - 3) - (k - 2)] 5 - 6k - 1 = 10k - 6 - 1k + 2 ? Stuck... Show steps please!

OpenStudy (nubeer):

u can simplify it more.. just add or subtract the terms with k and without k

OpenStudy (anonymous):

like 5-1 =4, 10k-1k =9k, -6+2 = -4 so 4 -6k = 9k-4 and then bring the ks on one side and the 4s of one side